OpenStudy (anonymous):

Solve. Check. Explain. 2x-5=y x-7=-y Addition/Elimination Method. (I'm a visual learner)

OpenStudy (ajprincess):

Add both the equations. Can u plz tell me vat u get @anonymous551?

OpenStudy (anonymous):

3x - 12y = 0

OpenStudy (ajprincess):

U r partly right. Plz check -12y

OpenStudy (anonymous):

3x - 12 = 0y ..? or 3x +12 = 0y

OpenStudy (ajprincess):

hmm it is 3x-12=0 2x-5=y-(1) x-7=-y-(2) (1)+(2)=>2x-5+x-7=y-y 3x-12=0 Nw using 3x-12=0 Can u find x @anonymous551?

OpenStudy (anonymous):

3x-12=0 +12 +12 3x = 12 --------- x = 4

OpenStudy (ajprincess):

ya perfect:) Now plug in that value of x in 2x-5=y to find y. Can u do it?

OpenStudy (anonymous):

2(4) - 5 = y 8 - 5 = y 3 = y

OpenStudy (ajprincess):

ya perfect:)

OpenStudy (anonymous):

NO. Sorry. 2x -5 = 3

OpenStudy (anonymous):

because 3 = Y

OpenStudy (ajprincess):

x=4 and y=3 is correct. Nw plug in these values in x-7=-y and check if Right hand side = left hand side

OpenStudy (anonymous):

Wait. One second..

OpenStudy (anonymous):

Oh.. Okay. 2x - 5 = y 2(4) - 5 = 3 8 - 5 = 3 True. x-7=-y (4) - 7 = -(3) -3 =-3 true.

OpenStudy (anonymous):

I though the equation was 2x - 12 for some reason.. Thanks for going through this w/ me.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

But you need to improve you explanation. If I didn't already know how to do this, I would have been so confused.

OpenStudy (ajprincess):

Welcome:) oh sorry. anyways thanx for the feedback:)